### v1

@@ Abstract
Storage systems account for a major portion of the total cost of ownership (TCO) of warehouse-scale computers, and thus have a major impact on the overall system’s efficiency. Machine learning (ML)-based methods for solving key problems in storage system efficiency, such as data placement, have shown significant promise. However, there are few known practical deployments of such methods. Studying this problem in the context of real-world hyperscale data center deployments at Google, we identify a number of challenges that we believe cause this lack of practical adoption. Specifically, prior work assumes a monolithic model that resides entirely within the storage layer, an unrealistic assumption in real-world data center deployments. We propose a cross-layer approach that moves ML out of the storage system and performs it in the application running on top of it, co-designed with a scheduling algorithm at the storage layer that consumes predictions from these application-level models. This approach combines small, interpretable models with a co-designed heuristic that adapts to different online environments. We build a proof-of-concept of this approach in a production distributed computation framework at Google. Evaluations in a test deployment and large-scale simulation studies using production traces show improvements of as much as 3.47$\times$ in TCO savings compared to state of the art baselines. We believe this work represents a significant step towards more practical ML-driven storage placement in warehouse-scale computers.
@@
Preliminary work.
@@ 1 Introduction
We posit that this is the key challenge behind existing proposed ML methods. Existing ML works mostly treat end-to-end data placement as one problem and assume a monolithic model deployed within the storage system Liu et al. (2020); Singh et al. (2022); Zhou & Maas (2021); Kaler & Toshniwal (2023). Such a model might be trained on file names or common application behavior Zhou & Maas (2021). While this approach works in simulation, it breaks the separation of concerns, which is problematic in real-world large-scale systems. For instance, changes to a major application that, e.g., affect file names would need to trigger a retraining of the model at the storage layer, which needs to roll out changes at a much lower velocity than the workloads.
Our design embraces the multi-layer characteristic of storage systems and presents a practical solution to these problems by combining cheap and interpretable ML models at the application layer with a custom algorithm that leverages their predictions at the storage layer. Instead of a single large ML model, we build smaller models for individual workloads, which produce hints that the storage layer can utilize to place the workloads’ data (Figure 2).
@@
Prior work has proposed imitation learning against such an oracle Liu et al. (2020). However, we find that this approach does not work in our deployment, since the model does not only need to make decisions for individual workloads but adapt to an environment that is changing due to external factors (e.g., varying load patterns and workloads arriving or leaving). We thus devise a cross-layer approach to tackle this adaptability problem. At the application layer, we analyze the data properties that contribute to the optimal placement. We then design a model to predict a ranking based on these properties, which is independent of online fluctuations of the environment and other applications. Finally, we design an adaptive algorithm at the storage layer to select the data to place on SSD based on the model predictions from all applications and system feedback.
We demonstrate our approach in the context of a real production data processing framework at Google. We show its practicality by developing a prototype of our approach and running it in a test deployment. We also perform an extensive simulation study based on real-world production traces from Google’s data centers. We show that our approach can lead to an additional 3.22% TCO savings, more than 3.47$\times$ the savings from the production baseline. We summarize our contributions as:
@@
  • We design and implement a novel cross-layer approach, combining ML and heuristics that can adapt across workloads and external factors in data centers.
@@ 2.3 Production Requirements and Limitations
Data centers accommodate a vast array of workloads with diverse behavior patterns. Employing a single model in the storage layer to jointly learn all workloads introduces a single point of failure, and requires the model to be large and complex, and thus expensive and difficult to interpret. Further, this approach requires all input features be reliably delivered to the storage system, when some of the most predictive features are workload-specific Zhou & Maas (2021). Finally, In hyperscale data centers, workloads exhibit significantly faster rates of change than the update cycles of storage systems. This causes a dilemma: 1) Rolling out the model with the storage system means it is stale by the time it reaches production; 2) updating the model independently of the storage system means that it is not tested as rigorously as the rest of the system, becoming its weakest link.
@@
Models can introduce non-trivial overheads. For example, prior work suggested using Transformers, which are known for their superior learning capabilities but can be computationally expensive, incurring prediction costs of approximately 99ms Zhou & Maas (2021). To balance performance and efficiency, we leverage gradient boosted trees as the model, providing a compromise between lightweight, low-performance models and powerful, expensive models.
@@ 3 Problem Formulation & Baselines
We evaluate resource savings from two perspectives. First, we measure the reduction in HDD I/O from moving jobs to SSDs. This frees up HDD I/O for operations that cannot (or should not) be moved off HDDs, such as accesses to cold data. We quantify this reduction using a metric called Total Cost of I/O (TCIO), where a value of 1.0 represents the amount of I/O that a standard HDD can sustain per second. Jobs running on SSDs have a TCIO of zero.
@@
$\mathrm{TCO^{\textsc{Dev}}}$ $=\mathrm{cost^{\textsc{Dev}}_{byte}}+\mathrm{cost^{\textsc{Dev}}_{network}}+\mathrm{cost^{\textsc{Dev}}_{server}}+\mathrm{cost^{\textsc{Dev}}_{specific}}$
@@
$\mathrm{cost^{SSD}_{server}}$ $=\mathrm{server\_cost\_rate^{SSD}}\cdot\mathrm{write\_throughput}$
@@
where _$*\_\mathrm{cost\_rate}$ denotes conversion rates to dollar cost; $\mathrm{cost^{\textsc{Dev}}_{byte}}$ denotes the cost of storing one byte for one second on a device; TCIO, size, and duration denote a job’s TCIO need, storage footprint, and duration (for example, if a job has a TCIO of 2, the job would need two HDDs to run). The $\mathrm{cost^{\textsc{Dev}}_{network}}$ is a value derived from the data center total network cost of transmitting data at the maximum throughput divided by the throughput per second. $\mathrm{cost^{HDD}_{server}}$ and $\mathrm{cost^{HDD}_{specific}}$ cover the cost of the servers and HDDs.
@@
Oracle: Optimal Solution Based on Solver. To better understand the headroom that is available if we achieve perfect data placement, we design an oracle. It is an upper bound on the best solution, but impossible to implement. The oracle policy is based on using an Integer Linear Programming (ILP) solver by assuming clairvoyant knowledge — that is, assuming we know the future access pattern. We formulate the placement problem as an ILP problem:
@@
FirstFit: Static Placement. Production systems commonly perform HDD/SSD tiering using FIFO or LRU-style heuristics Yang et al. (2023a); Eytan et al. (2020). We implement a representative instance of this approach. We try to place jobs on SSD in the order of their start times, checking jobs’ peak space usage and only placing jobs on SSD that fit the SSD availability. This optimizes TCIO when unlimited SSD is available but can significantly increase TCO when SSD capacity is limited or expensive.
Heuristic: Practical Adaptive Placement. Recently, heuristics that can dynamically adapt to workloads have been introduced, striking a balance between dynamically learning behavior and avoiding the practicality issues in Section 2.3. We emulate the state-of-the-art placement approach from Yang et al. (2022). It focuses on a slightly different problem (SSD read cache admission), but can be adapted for our placement task. We use this approach as a stand in for the closest practical approach to a learning-based baseline.
@@
ML Baseline: Lifetime Prediction-Based. We include another closely related ML approach Zhou & Maas (2021), which models storage problems in data centers as distribution prediction problems. We follow their SSD/HDD tiering case study to predict the mean ($\mu$) and standard deviation ($\sigma$) of a file lifetime. Files with a predicted lifetime ($\mu+\sigma$) shorter than the specified time-to-live (TTL) are admitted to SSD. To mitigate mispredictions, we evict any file residing in the SSD for longer than $\mu+\sigma$, as in their paper.
@@ 4 Hybrid Learning Approach
To address this and our other challenges (Section 2.3), we propose a cross-layer learning approach that uses the model only to predict a proxy for workload-specific characteristics and then co-designs a storage-level heuristic to turn the predictions into decisions for the current environment. Specifically, we want to design a proxy that correlates with a job’s TCO savings. The proxy allows predicted results to directly represent how a job’s placement contributes to end-to-end cost savings. We call this metric “importance” and train a categorical pointwise ranking model (category model) to learn the job’s importance ranking (Figure 3 right). Each category represents a particular ranking. A higher ranking indicates a more important job – placing it on SSD saves more cost. When making a placement decision, we query the model for an importance ranking category for a new job. We then run an adaptive category selection algorithm with dynamic feedback from the storage layer to decide which importance ranking categories to admit onto SSD.
@@ 4.1 Features
The features we choose (Table 2 in Appendix B) capture how these steps are executed (allocated resources, execution metadata). Execution metadata is formatted as strings that detail execution-related names, paths and targets. Key elements are separated by non-alphanumeric characters. We treat execution metadata as a sequence of substrings representing the key elements (Table 3 in Appendix B). Since executions may run periodically, we also include the weekday and hour of the day of a job’s start time. The allocated resource information represents resources assigned to the job by the cluster scheduler, before it starts execution. However, specific details regarding resource distribution, such as the assignment to SSD or HDD, are not determined at this stage. In addition, we also incorporate properties of previously completed jobs from the same user’s pipelines, including the past TCIO, job lifetime, and size. For a detailed description of these features, please refer to Appendix B.
@@ 4.2 Model Design
We use gradient boosted trees instead of neural networks (which are much more expensive and less interpretable) or lookup tables (which sacrifice accuracy). We build on the Yggdrasil Decision Forests framework Guillame-Bert et al. (2023). Our model is trained on the features in Section 4.1.
@@
$\textnormal{C}(x)=\begin{cases}0,&\text{if }x.m<0\\ k,&\text{if }x.n\in(\text{top }\frac{N-k}{N-1}*D,\text{top }\frac{N-k-1}{N-1}*D]\\ &\quad\text{and }x.m\geq 0\end{cases}$ (1)
@@ 4.3 Adaptive Category Selection Algorithm
We now discuss how our cross-layer design combines the learned category model and a heuristic-based algorithm for online data placement. As mentioned in Section 4, the model’s category prediction is independent of the SSD capacity. How to select the categories to place on SSD under different SSD capacities is non-trivial. A simple approach is to fix the admittable categories and always only place jobs predicted in these categories onto SSD. However, as shown in Figure 4, when we have larger SSD capacities, we want to dynamically admit more jobs compared to the smaller capacities case – that is, more categories.
@@
Algorithm Overview. Our algorithm makes the admission decisions based on real-time feedback regarding SSD utilization. When observing limited SSD capacity, we gradually decrease the number of categories to admit. Otherwise, we admit more categories. Since our category model predicts the “importance ranking” of jobs, admitting fewer categories naturally leads to admitting the most important jobs. Admitting more categories means that we broaden the admission set by adding less important but still cost saving jobs. We use a sliding category admission threshold to determine which predicted categories get placed on SSD.
SSD Usage Approximation. In cloud data centers, the actual SSD capacity varies among clusters of machines, which is also hard to directly approximate in practice. The criteria for determining whether an SSD is nearly full (i.e. can not fit more jobs) or not are influenced by workload patterns. Thus, we introduce a metric to unify the measurement of SSD capacity usage across clusters and workloads through job behavior observation. Given a sequence of jobs, $X=\langle x_{0},x_{1},\dots,x_{n}\rangle$, we introduce spillover TCIO percentage, $P_{\textsc{SpilloverTCIO}}{(X,t)}$, to measure the portion of all job $x_{i}$’s TCIO that is scheduled to be put onto SSD but ends up on HDD due to the fact that the SSD has already reached its full capacity at timestamp $t$ (notation in Table 1):
@@
Algorithm Design. We now introduce the adaptive category selection algorithm in Algorithm 1 with notation available in Table 1. In the algorithm, we keep track of an observation history $X_{h}$, which contains all the jobs starting within a fixed look back window, and calculate the SpilloverTCIO within the history, $h_{\textsc{SpilloverTCIO}}$. Then, we adaptively adjust the admission category threshold (ACT) based on the observed $P_{\textsc{SpilloverTCIO}}$ – if $P_{\textsc{SpilloverTCIO}}$ is larger than ACT, we increase the threshold to admit fewer categories. One issue of a dynamic control system of this kind is that ACT may change drastically. We combine two mechanisms to smooth the threshold change:
  • We use a spillover tolerance interval, $T_{\textsc{SpilloverTCIO}}$. When the observed SpilloverTCIO falls into the interval, we maintain ACT. If $P_{\textsc{SpilloverTCIO}}$ is below the lower bound of $P_{\textsc{SpilloverTCIO}}$, we decrease the threshold by $1$ (to avoid large adjustments). Conversely, for high $P_{\textsc{SpilloverTCIO}}$, we increase the ACT by $1$.
  • The threshold update is triggered by job arrival and the decision interval $t_{l}$ seconds instead of simply job coming to decrease the threshold change frequency.

### v2

@@ Abstract
Storage systems account for a major portion of the total cost of ownership (TCO) of warehouse-scale computers, and thus have a major impact on the overall system’s efficiency. Machine learning (ML)-based methods for solving key problems in storage system efficiency, such as data placement, have shown significant promise. However, there are few known practical deployments of such methods. Studying this problem in the context of real-world hyperscale data centers at Google, we identify a number of challenges that we believe cause this lack of practical adoption. Specifically, prior work assumes a monolithic model that resides entirely within the storage layer, an unrealistic assumption in real-world deployments with frequently changing workloads. To address this problem, we introduce a cross-layer approach where workloads instead “bring their own model”. This strategy moves ML out of the storage system and instead allows each workload to train its own lightweight model at the application layer, capturing the workload’s specific characteristics. These small, interpretable models generate predictions that guide a co-designed scheduling heuristic at the storage layer, enabling adaptation to diverse online environments. We build a proof-of-concept of this approach in a production distributed computation framework at Google. Evaluations in a test deployment and large-scale simulation studies using production traces show improvements of as much as 3.47$\times$ in TCO savings compared to state of the art baselines.
@@
Proceedings of the $\mathit{8}^{th}$ MLSys Conference, Santa Clara, CA, USA, 2025. Copyright 2025 by the author(s).
@@ 1 Introduction
We posit that this is the key challenge behind existing proposed ML methods. Existing ML works mostly treat end-to-end data placement as one problem and assume a monolithic model deployed within the storage system Liu et al. (2020); Singh et al. (2022); Zhou & Maas (2021); Kaler & Toshniwal (2023). Such a model might be trained on file names or common application behavior Zhou & Maas (2021). While this approach works in simulation, it breaks the separation of concerns, which is problematic in real-world large-scale systems. For instance, changes to a major application that affect file names would need to trigger a retraining of the model at the storage layer, which needs to roll out changes at a much lower velocity than the workloads.
We introduce a “Bring-Your-Own-Model” (BYOM) design that embraces the multi-layer characteristic of storage systems and presents a practical solution to these problems by combining cheap and interpretable ML models at the application layer with a custom algorithm that leverages their predictions at the storage layer. Instead of a single large ML model, we build smaller models for individual workloads, which produce hints that the storage layer can utilize to place the workloads’ data (Figure 2).
@@
Prior work has proposed imitation learning against such an oracle Liu et al. (2020). However, we find that this approach does not work in our deployment, since the model does not only need to make decisions for individual workloads but adapt to an environment that is changing due to external factors (e.g., varying load patterns and workloads arriving or leaving). Our BYOM design tackles this adaptability problem. At the application layer, we analyze the data properties that contribute to the optimal placement. We then design a model to predict a ranking based on these properties, which is independent of online fluctuations of the environment and other applications. Finally, we design an adaptive algorithm at the storage layer to select the data to place on SSD based on the model predictions from all applications and system feedback.
We instantiate this design in the context of a production data processing framework at Google. We show its practicality by developing a prototype of our approach and running it in a test deployment. We also perform an extensive simulation study based on real-world production traces from Google’s data centers. We show that our approach can lead to an additional 3.22% TCO savings, more than 3.47$\times$ the savings from the production baseline. We summarize our contributions as follows:
@@
  • We design and instantiate a novel cross-layer approach, combining ML and heuristics that can adapt across workloads and external factors in data centers.
@@ 2.3 Production Requirements and Limitations
Data centers accommodate a vast array of workloads with diverse behavior patterns. Employing a single model in the storage layer to jointly learn all workloads introduces a single point of failure, and requires the model to be large and complex, and thus expensive and difficult to interpret. Further, this approach requires all input features to be reliably delivered to the storage system, when some of the most predictive features are workload-specific Zhou & Maas (2021). Finally, In hyperscale data centers, workloads exhibit significantly faster rates of change than the update cycles of storage systems. This causes a dilemma: 1) Rolling out the model with the storage system means it is stale by the time it reaches production; 2) updating the model independently of the storage system means that it is not tested as rigorously as the rest of the system, thus becoming its weakest link.
@@
Models can introduce non-trivial overheads. For example, prior work suggested using Transformers, which are known for their superior learning capabilities but can be computationally expensive, incurring prediction latency costs of approximately 99ms Zhou & Maas (2021). To balance performance and efficiency, we leverage gradient boosted trees as the model, providing a compromise between lightweight, low-performance models and powerful, expensive models.
@@ 3 Problem Formulation & Baselines
We evaluate resource savings from two perspectives. First, we measure the reduction in HDD I/O from moving jobs to SSDs. This frees up HDD I/O for operations that cannot (or should not) be moved off HDDs, such as accesses to cold data. We quantify this reduction using a metric called Total Cost of I/O (TCIO), where a value of 1.0 represents the amount of I/O that a standard HDD can sustain per second. Jobs running on SSDs have a TCIO of zero. Our TCIO calculation accounts for caching effects, such as the DRAM cache present alongside HDDs in each server. In our system, I/Os that are served from cache do not reach the disks, and small write operations are grouped into 1 MiB chunks before reaching the disks. This ensures that TCIO reflects the true workload pressure on the disks.
@@
$\mathrm{TCO^{\textsc{Dev}}}$ $=\mathrm{cost^{\textsc{Dev}}_{byte}}+\mathrm{cost^{\textsc{Dev}}_{network}}+\mathrm{cost^{\textsc{Dev}}_{server}}$
[…]
@@
$\mathrm{cost^{SSD}_{server}}$ $=\mathrm{server\_cost\_rate^{SSD}}\cdot\mathrm{IO\_throughput^{SSD}}$
@@
where _$*\_\mathrm{cost\_rate}$ denotes conversion rates to dollar cost; $\mathrm{cost^{\textsc{Dev}}_{byte}}$ denotes the cost of storing one byte for one second on a device; TCIO, size, and duration denote a job’s TCIO need, storage footprint, and duration (for example, if a job has a TCIO of 2, the job would need two HDDs to run). The $\mathrm{cost^{\textsc{Dev}}_{network}}$ is a value derived from the data center total network cost of transmitting data at the aggregated throughput for the duration of the job. The network cost is largely constant and independent of the device, but we include it to avoid overestimating the impact of the other costs on the overall TCO. $\mathrm{cost^{HDD}_{server}}$ and $\mathrm{cost^{HDD}_{specific}}$ cover the cost of the servers and HDDs.
@@
3.1 Oracle: Optimal Solution Based on Solver.
To better understand the headroom that is available if we achieve perfect data placement, we design an oracle. It is an upper bound on the best solution, but impossible to implement. The oracle policy is based on using an Integer Linear Programming (ILP) solver and clairvoyant knowledge — that is, assuming we know the future access pattern. We formulate the placement problem as an ILP problem:
@@
3.2 FirstFit: Static Placement.
Production systems commonly perform HDD/SSD tiering using FIFO or LRU-style heuristics Yang et al. (2023a); Eytan et al. (2020). We implement a representative instance of this approach. We try to place jobs on SSD in the order of their start times, checking jobs’ peak space usage and only placing jobs on SSD that fit in the available SSD capacity. This optimizes TCIO when unlimited SSD is available but can significantly increase TCO when SSD capacity is limited or expensive.
3.3 Heuristic: Practical Adaptive Placement.
Recently, heuristics that can dynamically adapt to workloads have been introduced, striking a balance between dynamically learning workload behavior and avoiding the practicality issues in Section 2.3. We emulate the state-of-the-art placement approach from Yang et al. (2022). It focuses on a slightly different problem (SSD read cache admission), but can be adapted for our placement task. We use this approach as a stand in for the closest practical approach to a learning-based baseline.
@@
3.4 ML Baseline: Lifetime Prediction-Based.
We include another closely related ML approach Zhou & Maas (2021), which models storage problems in data centers as distribution prediction problems. We follow the paper’s SSD/HDD tiering case study to predict the mean ($\mu$) and standard deviation ($\sigma$) of a file lifetime. Files with a predicted lifetime ($\mu+\sigma$) shorter than the specified time-to-live (TTL) are admitted to SSD. To mitigate mispredictions, we evict any file residing in the SSD for longer than $\mu+\sigma$, as in the paper.
@@ 4 Hybrid Learning Approach
To address this and our other challenges (Section 2.3), we allow each workload to “bring their own model”. We propose a cross-layer learning approach that uses the model only to predict a proxy for workload-specific characteristics and then co-designs a storage-level heuristic that turns the predictions into decisions for the current environment. This design ensures that each workload’s I/O patterns is treated as independent, preventing one workload from impacting the I/O costs of others. Specifically, we want to design a proxy that correlates with a job’s TCO savings. The proxy allows predicted results to directly represent how a job’s placement contributes to end-to-end cost savings.
We call this proxy metric “importance” and train a categorical pointwise ranking model (category model) to learn the job’s importance ranking (Figure 3 right). Each category represents a particular level of importance. A higher ranking indicates a more important job – placing it on SSD saves more cost. When making a placement decision, we query the model for an importance ranking category for a new job. We then run an adaptive category selection algorithm with dynamic feedback from the storage layer to decide which importance ranking categories to admit onto SSD.
@@ 4.1 Features
The features we choose (Table 2 in Appendix B) capture how these steps are executed (allocated resources, execution metadata). Execution metadata is formatted as strings that detail execution-related names, paths and targets. Key elements are separated by non-alphanumeric characters. We treat execution metadata as a sequence of substrings representing the key elements (Table 3 in Appendix B). Since executions may run periodically, we also include the weekday and hour of the day of a job’s start time. Allocated resource information represents resources assigned to the job by the cluster scheduler, before it starts execution. However, specific details regarding resource distribution, such as the assignment to SSD or HDD, are not determined at this stage. In addition, we also incorporate properties of previously completed jobs from the same user’s pipelines, including the past TCIO, job lifetime, and size. For a detailed description of these features, please refer to Appendix B.
@@ 4.2 Model Design
We use gradient boosted trees instead of neural networks (which are much more expensive and less interpretable) or lookup tables (which sacrifice accuracy). We build on the Yggdrasil Decision Forests framework (YDF) Guillame-Bert et al. (2023). Our model is trained on the features in Section 4.1.
@@
$\textnormal{C}(x)=\begin{cases}0,&\text{if }x.m<0\\ k,&\text{if }x.n\in(\text{top }\frac{N-k}{N-1}*D,\text{top }\frac{N-k-1}{N-1}*D]\\ &\quad\text{and }x.m\geq 0\end{cases}$
@@ 4.3 Adaptive Category Selection Algorithm
We now discuss how our cross-layer design combines the learned category model and a heuristic-based algorithm for online data placement. As mentioned in Section 4, the model’s category prediction is independent of the SSD capacity. How to select the categories to place on SSD under varying SSD capacities is non-trivial. A simple approach is to fix the admittable categories and always only place jobs predicted in these categories onto SSD. However, as shown in Figure 4, when we have larger SSD capacities, we want to dynamically admit more jobs compared to the smaller capacities case – that is, more categories.
@@
Algorithm Overview. Our algorithm makes its admission decisions based on real-time feedback regarding SSD utilization. When observing limited SSD capacity, we gradually decrease the number of categories to admit. Otherwise, we admit more categories. Since our category model predicts the “importance ranking” of jobs, admitting fewer categories naturally leads to admitting the most important jobs. Admitting more categories means that we broaden the admission set by adding less important but still cost saving jobs. We use a sliding category admission threshold to determine which predicted categories get placed on SSD.
SSD Usage Approximation. In cloud data centers, the actual SSD capacity varies among clusters of machines, which is also hard to directly approximate in practice. The criteria for determining whether an SSD is nearly full (i.e. cannot fit more jobs) or not are influenced by workload patterns. Thus, we introduce a metric to unify the measurement of SSD capacity usage across clusters and workloads through job behavior observation. Given a sequence of jobs, $X=\langle x_{0},x_{1},\dots,x_{n}\rangle$, we define the spillover TCIO percentage, $P_{\textsc{SpilloverTCIO}}{(X,t)}$, to measure the portion of all job $x_{i}$’s TCIO that is scheduled to be put onto SSD but ends up on HDD due to the fact that the SSD has already reached its full capacity at timestamp $t$ (notation in Table 1):
@@
Algorithm Design. We now introduce the adaptive category selection algorithm in Algorithm 1 with notation available in Table 1. In the algorithm, we keep track of an observation history $X_{h}$, which contains all the jobs starting within a fixed look back window, and calculate the SpilloverTCIO within the history, $h_{\textsc{SpilloverTCIO}}$. Then, we adaptively adjust the admission category threshold (ACT) based on the observed $P_{\textsc{SpilloverTCIO}}$ – if $P_{\textsc{SpilloverTCIO}}$ is larger than ACT, we increase the threshold to admit fewer categories. One issue of a dynamic control system of this kind is that ACT may change drastically. We combine two mechanisms to smooth the ACT change:
  • We use a spillover tolerance range, $T_{\textsc{SpilloverTCIO}}$, within which the ACT remains unchanged. Avoiding large adjustments, if $P_{\textsc{SpilloverTCIO}}$ falls below the range lower bound, we decrease the threshold by 1; if $P_{\textsc{SpilloverTCIO}}$ exceeds the upper bound, we increase the ACT by 1.
  • This ACT update is triggered only at job arrivals and at fixed decision intervals of $t_{l}$ seconds, rather than on every job arrival. This reduces the frequency of threshold changes.